OpenStudy (anonymous):

Write an appropriate inverse variation equation if y = 9 when x = 3

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (zale101):

Y varies directly as x so y = kx. first, plug in y = 9 and x =3 to find k

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

9=k(3)

OpenStudy (anonymous):

so would it be y=3 over k?

OpenStudy (zale101):

oops k = 27

OpenStudy (anonymous):

or y=27 over k?

OpenStudy (zale101):

so y = k/x 9 = k/3 k = 27 y = 27/x